OpenStudy (anonymous):

How many moles of sulfuric acid are needed to produce 57.8 milliliters of water? Water has the density of 0.987 g/mL. Show all steps of your calculation as well as the final answer. NaOH + H2SO4 → Na2SO4 + H2O

OpenStudy (anonymous):

its pretty easy, follow these steps: balance equation: 2NaOH + H2SO4 --> Na2SO4 + 2H2O rho(density =d) = m/V -> m=d/V m=57,05g n=m/M=3,167mol n(H2SO4):n(H2O)=1:2 -> n(H2SO4) = 1/2 * n(H2O) = 0,5 * 3,167 mol = 1,583 mol